Prove the following using the following definition of O,Big-omega,Theta, small omega

Σki=1 ?i ?i = ?(nk )??? ? > 1.
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Let f(n)=a(1)*n+a(2)*n^2+.....a(k)*n^k

Since we know n^k>=n^c for any c<=k

So,

f(n)=a(1)*n+a(2)*n^2+.....a(k)*n^k<=(a(1)+a(2)+a(3)+.....a(k))*n^k

Let c=(a(1)+a(2)+a(3)+.....a(k))

So,

f(n)<=c*n^k

So, by deifinition of Big O

f(n)=O(n^k)

Kindly revert for any queries

Thanks.